CAVITY-ENHANCED, FREQUENCY-AGILE RAPID SCANNING (FARS) SPECTROSCOPY: EXPERI-MENTAL REALIZATIONS AND MEASUREMENT RESULTS

\underline{DAVID A. LONG}, GAR-WING TRUONG, ROGER VAN ZEE, DAVID F. PLUSQUELLIC, and JOSEPH T. HODGES, National Institute of Standards and Technology, 100 Bureau Drive, Gaithersburg, MD 20899, USA.

We present a series of experimental realizations of cavity-enhanced, frequency-agile rapid scanning (FARS) spectroscopy using distributed feedback diode lasers, external cavity diode lasers, and ultra-narrow linewidth fiber lasers. FARS offers a scanning rate which is limited only by the cavity response time itself as well as a microwave-level frequency axis. Finally, it allows for an absorption sensitivity which is one of the highest ever reported. These realizations offer a range of applications from low-cost field measurements of trace gases to laboratory-based metrology.